The sides of a right triangle are 8, 15 and 17 units. If each side is 1.5 times larger than original size, how many square units will be the area of new triangle?


1
Expert's answer
2021-05-17T17:08:15-0400

Area of right triangle is : 

"A = \frac{1}{2} × base × height"

base and Height is the small sides of right triangle and largest side is hypotenuse .

Let base = 8 units and height = 15 units

"Area= \frac{1}{2} × base × height"

"Area = \frac{1}{2} × 8 × 15"

"Area = 60 \space square\space units"

 

area of triangle after each side larger than 1.5 times the original sides 

"Let \space base = 1.5 × 8\space and\space Height = 1.5× 15"

"Now \space area = (\frac{1}{2}) × (1.5×8)×(1.5×15)"

"Area = 135 \space square\space units"

Area of new triangle = 135 square units .